If two solutions are mixed with the same kind of solute. If the conc. of Solution 1 has 2.979 M and has quantity of 9.164 L, and solution 2 has ___ and 6.271 L resulting to a Mixed conc. of 5.010 M.
1
Expert's answer
2021-10-01T08:30:03-0400

Solution:

M1 = Molarity of Solution 1 = 2.979 M

V1 = Volume of Solution 1 = 9.164 L

M2 = Molarity of Solution 2 = unknown

V2 = Volume of Solution 2 = 6.271 L

M = Molarity of mixture = 5.010 M


M = (M1V1 + M2V2) / (V1 + V2)

5.010 = (2.979 × 9.164 + M2 × 6.271) / (9.164 + 6.271)

5.010 = (27.2996 + M2 × 6.271) / (15.435)

5.010 × 15.435 = 27.2996 + M2 × 6.271

77.3294 - 27.2996 = M2 × 6.271

M2 × 6.271 = 50.0298

M2 = 7.978 M


Answer: Solution 2 has 7.978 M